# Hyperdimensional Probe: Decoding LLM Representations via Vector Symbolic Architectures

### v1

@@ 4.2 Input representations
The codebook defines the set of all input features; in our case, the contextually relevant concept, and is later used to construct and query VSA encodings. In our controlled setting, the codebook $\Phi$ (feature set) is constructed directly using all unique words included in the corpus, such as: $\texttt{mexico}\rightarrow\phi_{mexico}\in\Phi$, and $\texttt{krone}\rightarrow\phi_{krone}\in\Phi$. Thus, we create a matrix $\Phi\in\{-1,1\}^{n_{c}\times D}$, using the torch-hd library (Heddes et al., 2023), where $D$ is the VSA dimension and $n_{c}=2,996$ is the number of concepts/features. We set $D=4096$ as an adequate hidden dimension, given the cardinality of our codebook ($\approx 10^{3}$), which remains well below the theoretical capacity limit of the MAP-B architecture (Section˜3). The average pairwise cosine similarity of the concepts in the codebook is $0\pm 0.02$, confirming orthogonality (full distribution in Appendix˜J).
@@ 4.3 Processing LLM embeddings $F$
Our probing goal is to inspect the complete internal state of a language model prior to its next-token prediction, capturing all encoded concepts without assuming beforehand the type of relationship with its prediction. To this end, we examine the residual stream in the final token representation, focusing on the middle to last layers. Emerging evidence shows that transformers encode next-token information in the final token due to their autoregressive nature (Elhage et al., 2021; Olsson et al., 2022), refining it in later residual stream layers (Belrose et al., 2023; Hernandez et al., 2023). Specifically, for an autoregressive language model with $L$ hidden layers, we consider the embeddings (with size $d$) of the last token (“:”) in the latter half, for all $l\in[L/2,\dots,L]$, yielding a matrix in $\mathbb{R}^{L/2\times d}$.
@@
To reduce the computational cost of our approach, we lower the input dimensionality for our encoder by introducing two dimensionality-reduction steps: k-means clustering (Jain and Dubes, 1988), and sum pooling. Clustering reduces representation redundancy by grouping similar vector regions in LLM embedding space and computing centroids, accomplishing knowledge distillation. To determine the optimal range for $k$, we adopt the silhouette score (Rousseeuw, 1987). A trade-off between reduction, granularity, and model variability emerges with 3–7 clusters (Section˜O.3). We set $k=5$ to maintain the essential data structure while supporting effective dimensionality reduction. [^2] Section O.4 shows that the clusters consistently group adjacent layers. We then apply sum pooling, which consists of summing all centroid embeddings; [^3] Preliminary evidence suggests that directly summing all layers (up to 32) results in a noisier representation. merging group representatives (k-dimensional matrix) into a vector exploiting the additivity property of LLM embeddings demonstrated in previous work (Bronzini et al., 2024). For example, these reduction steps allows us to downsize the probed embedding space of OLMo-2:
@@ 4.4 Neural VSA encoder $T$
We valide our methodology on embeddings from popular open-weight LLMs available on the Hugging Face platform with 355M-109B parameters, experimenting with different embedding sizes and layer counts. In particular, we test the latest Meta AI’s Llama 4 Scout, (AI, 2025) a multi-modal mixture of 16 experts (MoE), Llama 3.1-8B (Grattafiori et al., 2024), Microsoft’s Phi-4 (Abdin et al., 2024), EleutherAI’s Pythia-1.4b (Biderman et al., 2023), AllenAI’s OLMo-2-32B (OLMo et al., 2024), and OpenAI’s legacy GPT-2-medium (Radford et al., 2019).
@@
The LLM-VSA dataset uses a random 70-15-15 split of $\mathcal{S}$ for training, validation, and test sets. Since our setting can be interpreted both as a vector-based regression task and a multi-label classification problem, [^5] VSA encodings can be viewed as vectors with $D$ distinct labels, each assuming one of two possible values. we evaluate our approach using two distinct metrics: cosine similarity and multi-label binary accuracy. For binary accuracy, targets and predictions are binarized based on sign. First, evaluating the cosine similarity between the predicted and target VSA encodings yields a test-set average score of $0.89$ (best LLM in Appendix˜D, Llama 3.1-8B), indicating strong numerical alignment between our encoder’s outputs and the target encodings. Second, we obtain an average binary accuracy of $0.94$, which indicates robust classification accuracy after polarizing the predictions with the sign function. This means that on average, the VSA encodings produced by our trained model deviated from the targets by only $6\%$ of the vector elements, a negligible error given VSA’s large tolerance to noise. All tested models exhibit consistent performance; [^6] Appendix D reports the training performance of our neural VSA encoder $\mathcal{M}$ for all of the six models. layer count has no effect, whereas reducing the embedding dimension is found to be slightly detrimental. This empirical evidence supports the effectiveness of our proposed methodology and the hypothesis that LLM embeddings can be represented using fully distributed encodings such as MAP-B in VSAs.
[…]
@@ 4.5 Probing VSA encodings II
When probing an encoding ($y_{s}$ in Equation˜4), we pick in-context concepts $(\phi_{\textrm{denmark}}$, $\phi_{\textrm{krone}}$, and $\phi_{\textrm{mexico}})$, and their combinations, as candidates for unbinding. The best candidate was chosen by benchmarking the resulting concept after unbinding, against the in-context and target concepts through cosine similarity. If no relevant match was found ($sim<0.1$), no operation was applied. In the experiments reported in Section˜5, 80% of unbinding operations, averaged across all models, relied on the key of the target pair. In contrast, no operation was applied in 12% of the cases. Appendix˜E shows the proportions of other candidates and highlights the variations among models.
5 Experiments and results
This section presents insights into experiments with our trained encoders. We first outline the experimental setup in Section˜5.1. We then report findings on LLM performance and concept extraction using our hyperdimensional probe in Section˜5.2. Appendix˜G contrasts our results with DLA, showing inferior probing capabilities, likely due to its reliance on the next-token representation, and thus surface-level features. Lastly, Section˜5.3 extends our approach to the question-answering task.
@@ 5.1 Experimental setup
We test our trained neural VSA encoders $\mathcal{M}$ on the set of textual inputs formatted using the verbose template ($\mathcal{\bar{S}}$, Section˜4.1). Thus, we validate our methodology using inputs with syntactic structures that differ from those seen during the training stage. Therefore, we perform information decoding from the vector representation of a different token, shifting from the colon token of the training template (Equation˜1) to the token to. This aims to further mitigate confounding effects from probe’s task performance in relation to information decoding.
@@
Our experimental evaluation has a two-fold objective (Equation˜4): we assess the performance of LLMs in the next-token prediction, and our VSA-based probing method for retrieving targets from their latent representations using precision@k. We measure the LLM’s performance via: binary precision on the next-token prediction against the target word; softmax score of the most likely next token and the target one; and rank of the target token on the ordered softmax scores. We compute the precision of LLM predictions by considering the most likely next token based on softmax scores (next-token@1), and the top-5 most likely ones (next-token@5). To address scenarios where the token generated by the LLM includes the initial part of the target word due to tokenization, we introduce a value of 0.5 in the LLM’s precision metrics (see also Appendix˜A). For example, this value is assigned if the model predicts the next token as ack for the target word acknowledge. To evaluate the performance of our VSA-based probing approach, we assess the binary precision of retrieving the target VSA concept from LLM latent representations via probing@1, and probing@5.
5.2 Extracting next-token concepts
Figure˜4 shows VSA-based probing for target concept retrieval, and LLM performance to complete analogies with targets. Appendix˜F displays the same results in a table, with statistical variability and two control tests.
High variability in LLM performance.
@@ 5.2 Extracting next-token concepts
In an unexpected contrast, the largest model evaluated (109B; Llama 4, Scout) exhibited the lowest precision@1 in the next-token prediction task (8%, Figure˜4), even underperforming the legacy GPT-2. Yet, its next-token@5 was comparable to others (still the lowest), but ranked among the best in probing@1. Strong probing performance suggests the final state encodes the target concept, but the model often fails to output it. This might be caused by exogenous (e.g., prompt design) and endogenous factors (e.g., tokenization). As shown in Section˜F.3, the model frequently predicted a space instead of the correct word, which still often appeared in its top five predictions. This emphasizes variability introduced by tokenization and prompt design, which might have greater impact on token-based probing methods such as DLA.
Figure 4: LLM performance in completing the analogies with target words (left), and the effectiveness of our decoding method in extracting the targets from LLM latent representations (right).
@@
Key $\mid$ Target 60.0 66.9 85.4 84.8 80.1 79.0 76.0 ± 9.4
NONE 21.9 16.7 6.9 7.6 8.5 11.5 12.2 ± 5.4
Key 4.5 6.1 0.6 1.0 1.8 3.4 2.9 ± 2.0
Example 5.8 2.4 1.5 1.3 2.0 0.8 2.3 ± 1.6
Context $\mid$ Target 1.1 1.9 1.5 1.2 4.4 2.4 2.1 ± 1.1
Key $\mid$ Key Values 1.3 1.4 1.8 1.5 1.1 0.8 1.3 ± 0.3
Out-of-context 1.6 1.2 0.5 0.7 0.8 1.1 1.0 ± 0.4
Example Value $\mid$ Key Values 1.5 1.3 0.3 0.0 0.2 0.1 0.6 ± 0.6
Key Values $\mid$ Target 0.4 0.1 0.3 0.4 0.1 0.1 0.2 ± 0.1
Target 0.1 0.1 0.1 0.2 0.1 0.1 0.1 ± 0.0
VSA probing exposes varying conceptual richness.
Regarding the concepts extracted by our hyperdimensional probe, we achieve an average probing@1 across all models equal to 83% (right side of Figure˜4, Appendix˜F), extracting the target concept with its key for most cases (60% for GPT-2, 85% for Llama 3.1, Table˜1). Notably, GPT-2 shows the highest percentage of cases where no concepts are extracted (22%) and only the concepts from the in-context example (6%), ranking second in extracting only the keys of the target concept (5%). This underscores its struggle with the NLP task of completing analogies, even when it appears to grasp the context. On the other hand, OLMo-2 has the highest proportion of instances in which our probing approach retrieves the target concept alongside all in-context concepts (Context $\mid$ Target, 4%), indicating its richer representation in its final state for both the input context and next word. This latent richness is then reflected in its performance on next-token prediction, achieving the highest next token@1 equal to 48% (Figure˜4). In cases where the target word was not among the top five predictions of Llama 4, nearly 50% of the instances (Section˜L.1; 28% for OLMo-2 in Section˜L.2), [^7] We report the target word is absent from next token@k, including also tokens that represent its beginning. our probing method successfully extracted the target concept and its associated key in 70% of instances, while no concept was retrieved in 18% of cases (26% for OLMo-2). Although the first outcome supports previous observations, the absence of extracted concepts merits a more granular analysis across analogy categories (Table˜12 in Appendix˜M). Our probe most frequently encounters conceptually-empty representations in mathematical analogies (88%, Section˜F.2, also for OLMo-2 comparison), followed by semantic hierarchies (39%). Factual and morphological analogies show much lower rates, at 5.5% and 1.1%, respectively. As elaborated in Section˜F.2, these differences likely stem from the type of reasoning involved: linguistic analogies depend on syntactical patterns, factual and semantic relations on key–value associations, and hierarchies or mathematical analogies on abstract inference.
5.3 From input-completion tasks to question-answering
To further validate our approach in real-world scenarios and beyond the controlled analogy task explored previously, we apply our methodology to a question answering task, using the popular SQuAD dataset (Rajpurkar et al., 2016).

### v2

@@ 4.2 Training examples
The codebook defines the set of all input features; in our case, the contextually relevant concept, and is later used to construct and query VSA encodings. In our controlled setting, the codebook $\Phi$ (feature set) is constructed directly using all unique words included in the corpus, such as: $\texttt{mexico}\rightarrow\phi_{mexico}\in\Phi$, and $\texttt{krone}\rightarrow\phi_{krone}\in\Phi$. Thus, we create a matrix $\Phi\in\{-1,1\}^{n_{c}\times D}$, using the torch-hd library (Heddes et al.,, 2023), where $D$ is the VSA dimension and $n_{c}=2,996$ is the number of concepts/features. We set $D=4096$ as an adequate hidden dimension, given the cardinality of our codebook ($\approx 10^{3}$), which remains well below the theoretical capacity limit of the MAP-B architecture (Section˜3). The average pairwise cosine similarity of the concepts in the codebook is $0\pm 0.02$, confirming orthogonality (full distribution in Appendix˜J).
@@ 4.3 Processing neural embeddings $F$
Our probing goal is to inspect the complete internal state of a language model prior to its textual output, capturing all encoded concepts without assuming their relation to the output. To this end, we examine the residual stream in the final token representation, focusing on the middle to last layers. Emerging evidence shows that transformers encode next-token information in the final token due to their autoregressive nature (Elhage et al.,, 2021; Olsson et al.,, 2022), refining it in later residual stream layers (Belrose et al.,, 2023; Hernandez et al.,, 2023). Specifically, for an autoregressive language model with $L$ hidden layers, we consider the embeddings (with size $d$) of the last token (“:”) in the latter half, for all $l\in[L/2,\dots,L]$, yielding a matrix in $\mathbb{R}^{L/2\times d}$.
@@
To reduce the computational cost of our approach, we lower the input dimensionality for our encoder by introducing two dimensionality-reduction steps: k-means clustering (Jain and Dubes,, 1988), and sum pooling. Clustering reduces representation redundancy by grouping similar vector regions in LLM embedding space and computing centroids, accomplishing knowledge distillation. To determine the optimal range for $k$, we adopt the silhouette score (Rousseeuw,, 1987). A trade-off between reduction, granularity, and model variability emerges with 3–7 clusters (Section˜O.3). We set $k=5$ to maintain the essential data structure while supporting effective dimensionality reduction. [^2] Section O.4 shows that the clusters consistently group adjacent layers. We then apply sum pooling, which consists of summing all centroid embeddings; [^3] Preliminary evidence suggests that directly summing all layers (up to 33) results in a noisier representation. merging group representatives (k-dimensional matrix) into a vector exploiting the additivity property of LLM embeddings demonstrated in previous work (Bronzini et al.,, 2024). For example, these reduction steps allows us to downsize the probed embedding space of OLMo-2:
@@ 4.4 Neural VSA encoder $T$
We validate our methodology on embeddings from popular open-weight LLMs available on Hugging Face with 355M-109B parameters, experimenting with different embedding sizes and layer counts. In particular, we test the latest Meta AI’s Llama 4 Scout, (Meta AI,, 2025) a multi-modal mixture of 16 experts (MoE), Llama 3.1-8B (Grattafiori et al.,, 2024), Microsoft’s Phi-4 (Abdin et al.,, 2024), EleutherAI’s Pythia-1.4b (Biderman et al.,, 2023), AllenAI’s OLMo-2-32B (OLMo et al.,, 2024), and OpenAI’s legacy GPT-2-medium (Radford et al.,, 2019).
@@
The LLM-VSA dataset uses a random 70-15-15 split of $\mathcal{S}$ for training, validation, and test sets. Since our setting can be interpreted both as a vector-based regression task and a multi-label classification problem, [^5] VSA encodings can be viewed as vectors with $D$ distinct labels, each assuming one of two possible values. we evaluate our approach using two distinct metrics: cosine similarity and multi-label binary accuracy. For binary accuracy, targets and predictions are binarized based on sign. First, evaluating the cosine similarity between the predicted and target VSA encodings yields a test-set average score of $0.89$ (best LLM in Appendix˜D, Llama 3.1-8B), indicating strong numerical alignment between our encoder’s outputs and the target encodings. Second, we obtain an average binary accuracy of $0.94$, which indicates robust classification accuracy after polarizing the predictions with the sign function. This means that on average, the VSA encodings produced by our trained model deviated from the targets by only $6\%$ of the vector elements, a negligible error given VSA’s large tolerance to noise. All tested models exhibit consistent performance; [^6] Appendix D reports the training performance of our neural VSA encoder $\mathcal{M}$ for all of the six models. layer count has no effect, whereas reducing the embedding dimension is found to be slightly detrimental. These results support our empirical hypothesis that VSA-based representations (MAP-B) and hypervector algebra can faithfully capture the latent features encoded in the neural embedding of LLMs (RQ7).
@@ 4.5 Probing VSA encodings II
When probing an encoding ($y_{s}$ in Equation˜4), we pick in-context concepts $(\phi_{\textrm{denmark}}$, $\phi_{\textrm{krone}}$, and $\phi_{\textrm{mexico}})$, and their combinations, as candidates for unbinding. The best candidate was chosen by benchmarking the resulting concept after unbinding, against the in-context and target concepts through cosine similarity with a threshold for concept detection equal to $0.1$. Empirical tests show relevant concepts exceeded this low threshold, while noisy ones remain below, likely due to VSA’s noise tolerance. If no concepts were detected, unbinding was skipped. In the experiments reported in Section˜5, 80% of unbinding operations, averaged across all models, relied on the key of the target pair. In contrast, no operation was applied in 12% of the cases. Appendix˜E shows the proportions of other candidates and highlights the variations among models. Meanwhile, the unrelated baseline in Appendix˜F shows results of unbinding with concept candidates that do not relate to the input.
Figure 4: LLM performance on analogy completion with targets (left) and our decoding method’s ability to extract targets from LLM embeddings (right), with logit-based results shown in the middle.
5 Experiments on input-completion tasks
This section explores the model’s internal state when completing analogy-style inputs (RQ1–RQ5). We begin by outlining the analogy-oriented setup in Section˜5.1. Next, we present our findings on concept extraction using the proposed probe in Section˜5.2. Lastly, Section˜5.3 compares these results with traditional logit-based investigation, which exhibits more superficial probing capabilities, likely because it depends on the model’s vocabulary space and, thus, token-level latent features.
@@ 5.1 Experimental setup
We test our trained neural VSA encoders $\mathcal{M}$ on the set of textual inputs formatted using the verbose template (Section˜4.1). Thus, we validate our methodology using inputs with syntactic structures that differ from those seen during the training stage ($\mathcal{\bar{S}}\neq\mathcal{S}$). Therefore, we also perform information decoding from the vector representation of a different token, shifting from the colon token of the training examples (Equation˜1) to the token to in $\mathcal{\bar{S}}$. This further mitigates confounding effects caused by probe-induced learning, more pronounced in conventional supervised probes.
@@
Our experimental evaluation has a two-fold objective (Equation˜4): we assess the performance of LLMs in the next-token prediction, and our VSA-based probing method for retrieving targets from their latent representations using precision@k. We measure the LLM’s performance via: binary precision on the next-token prediction against the target word; softmax score of the most likely next token and the target one; and rank of the target token on the ordered softmax scores. Downstream performance of LLMs is measured using the most likely next token (next-token@1) and the top-5 tokens (next-token@5), capturing uncertainty in the model’s softmax distribution. To evaluate the performance of our VSA-based probing approach, we assess the binary precision of retrieving the target VSA concept from LLM latent representations via probing@1, and probing@5.
5.2 Concept extraction
Figure˜4 reports VSA-based target concept retrieval, LLM downstream performance, and logit-based comparison. Appendix˜F tabulates the same results with variability and two validation tests. Table˜1 presents the final results of VSA probing, with each row showing the unique concept sets produced by the extraction procedure of Section˜4.5. Each set includes the concepts used for unbinding and those identified afterward; $\phi_{\text{mexico}}\mapsto$ Key, and $\phi_{\text{peso}}\mapsto$ Target for the example within Equation˜4.
High variability in LLMs’ next-token predictions
@@ 5.2 Concept extraction
. In an unexpected contrast, the largest model evaluated (109B; Llama 4, Scout) exhibited the lowest precision@1 in the next-token prediction task (8%, Figure˜4), even underperforming the legacy GPT-2. Yet, its next-token@5 was comparable to others (still the lowest), but ranked among the best in probing@1. Strong probing performance suggests the final state encodes the target concept, but the model often fails to output it. This might be caused by exogenous (e.g., prompt design) and endogenous factors (e.g., tokenization). As shown in Section˜F.3, the model frequently predicted a space instead of the correct word, which still often appeared in its top five predictions. This emphasizes variability introduced by prompt design and tokenization, which might have greater impact on logit-based methods such as DLA.
VSA probing exposes varying conceptual richness (RQ1–RQ5).
Regarding the concepts extracted by our hyperdimensional probe, we achieve an average probing@1 across all models equal to 83% (right side of Figure˜4, Appendix˜F), extracting the target concept with its key for most cases (60% for GPT-2, 85% for Llama 3.1, Table˜1). Notably, GPT-2 shows the highest proportion of cases with no concepts extracted (22%) followed by cases extracting only the in-context example concepts (6%), and ranks second in extracting only the target concept keys (5%). Thus, its internal states largely fall back on input-related concepts, reflecting limited understanding needed to complete the analogy-style input correctly. On the other hand, OLMo-2 has the highest proportion of instances in which our probing approach retrieves the target concept alongside all in-context concepts (Context $\mid$ Target, 4%), indicating its richer representation in its final state for both the input context and next word. This latent richness is then reflected in its performance on next-token prediction, achieving the highest next token@1 equal to 48% (Figure˜4). In cases where the target word was not among the top five predictions of Llama 4, nearly 50% of the instances (Section˜L.1; 28% for OLMo-2 in Section˜L.2), our probing method successfully extracted the target concept and its associated key in 70% of instances, while no concept was retrieved in 18% of cases (26% for OLMo-2). Although the first outcome supports previous observations, the absence of extracted concepts merits a more granular analysis across analogy categories (RQ4; see also Table˜12 in Appendix˜M). Our probe most frequently encounters conceptually-empty representations in mathematical analogies (88%, Section˜F.2, also for OLMo-2 comparison), followed by semantic hierarchies (39%). Factual and morphological analogies show much lower rates, at 5.5% and 1.1%, respectively. As elaborated in Section˜F.2, these differences likely stem from the type of reasoning involved: linguistic analogies depend on syntactical patterns, factual and semantic relations on key–value associations, and hierarchies or mathematical analogies on abstract inference.
@@
Key $\mid$ Target 60.0 66.9 85.4 84.8 80.1 79.0 76.0 $\pm$ 9.4
NONE 21.9 16.7 6.9 7.6 8.5 11.5 12.2 $\pm$ 5.4
Key 4.5 6.1 0.6 1.0 1.8 3.4 2.9 $\pm$ 2.0
Example 5.8 2.4 1.5 1.3 2.0 0.8 2.3 $\pm$ 1.6
Context $\mid$ Target 1.1 1.9 1.5 1.2 4.4 2.4 2.1 $\pm$ 1.1
Key $\mid$ Key Values 1.3 1.4 1.8 1.5 1.1 0.8 1.3 $\pm$ 0.3
Out-of-context 1.6 1.2 0.5 0.7 0.8 1.1 1.0 $\pm$ 0.4
Example Value $\mid$ Key Values 1.5 1.3 0.3 0.0 0.2 0.1 0.6 $\pm$ 0.6
Key Values $\mid$ Target 0.4 0.1 0.3 0.4 0.1 0.1 0.2 $\pm$ 0.1
Target 0.1 0.1 0.1 0.2 0.1 0.1 0.1 $\pm$ 0.0
GPT-2 Pythia Llama 3.1 Phi-4 OLMo-2 Llama 4, Scout AVERAGE
NONE: no concepts from DLA (%; $\mathcal{D}\subset\bar{\mathcal{S}}$) 33.9 32.8 47.4 33.1 14.6 15.4 29.5 $\pm$ 11.4
Concepts extracted from $\mathcal{D}$ by VSA (%)
Key $\mid$ Target 53.5 56.5 76.6 70.5 44.5 42.6 57.4 $\pm$ 12.5
NONE 26.8 24.3 13.7 18.6 41.0 43.2 27.9 $\pm$ 10.9
Example 6.8 3.0 2.1 2.1 3.4 2.0 3.2 $\pm$ 1.7
Key 5.2 7.3 0.7 0.9 1.6 3.0 3.1 $\pm$ 2.4
Out-of-context 2.0 1.8 1.1 1.9 3.8 4.5 2.5 $\pm$ 1.2
Key $\mid$ Pair Values 1.8 2.1 2.5 3.2 2.4 1.8 2.3 $\pm$ 0.5
Context $\mid$ Target 0.7 1.2 1.1 0.7 1.9 1.4 1.2 $\pm$ 0.4
Target 0.1 0.1 0.2 0.0 0.1 0.0 0.1 $\pm$ 0.1
5.3 Comparison with Logit Attribution
For validation, we compare with conventional output distribution inspection. Specifically, we apply DLA (Logit Lens; Section˜2) to all models using $\bar{S}$ to highlight the limitations of token-aligned logit–based analysis. Appendix˜G provides details on our comparison choices and further analysis.
We use fuzzy token-to-concept matching with our concept set (e.g., “pes" $\mapsto$ peso), and consider projected next-token predictions from the model’s middle to last layers (Section˜G.3) of the last token, as VSA probing. DLA produces no concepts in nearly 30% of analogies on average ($\mathcal{D}$ in Table˜2; +17% compared to VSA in Table˜1). It also exhibits lower performance and greater variability in target concept retrieval, with a similarly large drop in comparison to VSA-based precision@1 (Figure˜4). In instances without concepts from DLA ($\mathcal{D}$ in Table˜2), our VSA probing extracts, on average, the key-target pair in 57% of all analogies, while returning none for 28%. For instance, for the analogy king is to queen as son is to $\mapsto$ daughter, using OLMo-2, our probe extracts the key-target concepts (son and daughter), while DLA produces no concepts. The model predicts the next token prediction as ? with a softmax score of 0.06, followed by father (0.05); the target word has a rank of 37. Since conventional output inspections operate within the model’s vocabulary space, token-level logit analyses often produce shallower findings and show greater variability caused likely by exogenous factors, such as prompt design, or endogenous ones.
6 From input-completion tasks to text generation
Lastly, this section examines the model’s internal state during text generation (RQ6). We apply our methodology (Section˜4) to a question-answering scenario, which represents a more realistic setting than the previous controlled testbed, using the popular SQuAD dataset (Rajpurkar et al.,, 2016).